OpenStudy (anonymous):

Prove that\[\left(m^{(p-1)(q-1)}\right)^km\equiv m\text{ }(\text{mod }pq),\]where \(m

OpenStudy (across):

Assuming that \(m\) is a positive integer, if \(\gcd(m,pq)\neq1\), then \(m\) is a multiple of either \(p\) or \(q\). Without loss of generality, let \(m=p\). Then we have\[\left(p^{(p-1)(q-1)}\right)^kp\equiv p\text{ }(\text{mod }pq)\]and\[\left(p^{(p-1)(q-1)}\right)^k\equiv 1\text{ }(\text{mod }q).\]However,\[\left(p^{k(p-1)}\right)^{q-1}\equiv 1\text{ }(\text{mod }q),\]which is true by Fermat's little theorem. Can you wrap it up from here?

OpenStudy (anonymous):

yes. thank you very much